INFRARED SPECTROSCOPY OF CARBON- AND CARBON-SILICON CLUSTERS

J. KRIEG, V. LUTTER, I. GOTTBEHÜT, T. F. GIESEN, S. SCHLEMMER, and \underline{S. THORWIRTH}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

Many of the molecules found in space are carbonaceous, that is, they have a carbon backbone in their structure. In addition, many of these molecules carry heteroatoms such as nitrogen and oxygen and also second row elements such as silicon. To date, four silicon-carbon molecules SiC_{n} (n = 1 − 4) have been detected in space and several more by high-resolution spectroscopic techniques in the laboratory. Owing to their symmetry, many clusters of the form SiC_{n}Si (and linear C_{n} chains) are non-polar and hence have no pure rotational spectrum. In an effort to obtain the gas-phase spectra of these clusters in the infrared, we have started a dedicated laboratory program employing diode laser techniques and more recently an optical parametric oscillator-based spectrometer operating at 5 microns, where many carbon- and carbon-silicon chains are expected to exhibit strong infrared-active vibrational modes. Results from new observations of the previously studied Si_{2}C_{3} and C_{6} clusters ^{a}^{,}^{b} will be reported.

aA. Van Orden, T. F. Giesen, R. A. Provencal, H. J. Hwang, and R. J. Saykally, J. Chem. Phys. 101, 10237 (1994).

bH. J. Hwang, A. van Orden, K. Tanaka, E. W. Kuo, J. R. Heath, and R. J. Saykally, Mol. Phys. 79, 769 (1993).